if a,b,c,d,e,f are non negative real numbers and a+b+c+d+e+f=1 then maximum value of ab+bc+cd+de+ef
1
Expert's answer
2012-08-21T09:24:50-0400
(a+b+c+d+e+f)^2=1
A^2+b^2+c^2+d^2+e^2+f^2+2(ab+bc+cd+de+ef) = 1
(ab+bc+cd+de+ef)<1/2
Ans 1/2